OpenStudy (anonymous):

help please!! The graph of which of the following will be perpendicular to the graph of y = 2x – 4? A. 2x + y = 4 B. 4x – 2y = 4 C. y = –2x + 8 D. 2x + 4y = 1

OpenStudy (anonymous):

There are two things you should keep in mind. 1. Slope of line ax+by-c=0 is -a/b. 2. Two lines will be perpendicular if the product of their slopes is -1. Here, Slope of given line y=2x-4 is 2. So, the line which is perpendicular to it should have a slope of -1/2. Now, Slope of line 2x+y=4 is -2. Slope of line 4x-2y=4 is 2. Slope of line y=-2x+8 is -2. Slope of line 2x+4y=1 is -1/2. So, Right ans. is D. 2x+4y=1.